# Spectral Compressive Imaging via Chromaticity-Intensity Decomposition

### v1

@@ 2.3 Hybrid Spatial-Spectral Transformer
We adopt a dual-branch design: the spatial attention branch captures intra-image textures through a Swin Transformer in Encoder, while the spectral attention branch is tailored to exploit sparse and locally correlated spectral features using a TopK spectral attention mechanism in Decoder. This asymmetric design is inspired by zhang2024dual and motivated by experimental verification, where the asymmetric design has better reconstruction results.
@@
where $\mathcal{R}$ denotes the set of TopK ratios. This fusion across multiple sparsity levels enables the network to capture both dominant and complementary spectral correlations, further improving spectral detail preservation. and reshaped back to the original window structure. After processing all windows, the outputs are stitched to reconstruct the full feature map. This spectral TopK attention not only reduces the computational burden from dense $O(C^{2})$ to $O(KC)$, but also explicitly captures the structured sparsity observed in reflectance spectra—where only a few wavelengths contribute significantly. Empirically, we find that this strategy enhances spectral sharpness and suppresses irrelevant cross-band mixing.
@@
Inspired by recent advances in Vision Transformers liu2021swin , we employ the Swin Transformer as our spatial modeling backbone. The input feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ is divided into non-overlapping windows of size $M\times M$. Within each window, we perform multi-head self-attention (MSA) by computing
@@ 3 Experiments
Simulation Dataset. We adopt two widely used hyperspectral datasets: CAVE yasuma2010generalized and KAIST choi2017high . The CAVE dataset contains 32 hyperspectral images with a spatial resolution of $512\times 512$ and 31 spectral bands. The KAIST dataset provides 30 high-resolution hyperspectral scenes of size $2704\times 3376$ with 31 bands. We obtain ground-truth multi-spectral chromaticity and intensity using chromaticity-intensity decomposition. Following prior works meng2020end ; miao2019net ; cai2022mask , we use all CAVE images for training and select 10 scenes from KAIST for evaluation.
Implementation Details. Our model is implemented in PyTorch and trained using the Adam optimizer kingma2014adam for 300 epochs. The initial learning rate is set to $4\times 10^{-4}$ and updated using a cosine annealing schedule. We employ the $\ell_{2}$ loss between the reconstructed chromaticity and ground-truth chromaticity as the objective function. For training, we randomly extract 3D hyperspectral patches from each scene. For simulated data, the patch size is $256\times 256\times 28$, for real-world data, we use patches of size $350\times 260\times 26$. In simulation experiments, the forward imaging model is configured with a dispersion shift step $d=2$, directing dispersion along the horizontal axis (rightward). In real-world scenarios, we assume a vertical dispersion direction and set $d=1$, consistent with the dual-camera hardware setup. All experiments are conducted in Nvidia A40 GPU.
@@
Comparing Methods. We compared the HSIs reconstruction performance of our CIDNet with other 7 SOTA methods, including DeSCIliu2018rank , GAP-Netmeng2020gap , MST-Lcai2022mask , DAUHST-9stgcai2022degradation , SSR-9stgzhang2024improving and two dual-camera CASSI algorithm: PIDSchen2023prior and In2SET-9stgwang2024in2set . PIDS is compared with RGB image as guidance(in original paper). We compared the spectral reconstruction quality in two aspects, the original HSIs and the chromaticity. Note that our method reconstruct the chromaticity while other baseline algorithms reconstruct the HSIs. Therefore for fair comparison, the chromaticity the chromaticity is computed based on the chromaticity-intensity decomposition after HSIs reconstruction. We multiply the known intensity with the recovered chromaticity to obtain its HSIs in our method. The reconstruction quality of hyperspectral images are assessed using peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) metrics.
@@ 3.1 Quantitative Results
As shown in Tab. 1, we compare the PSNR and SSIM of HSIs with SOTA methods. our proposed CIDNet excels in 8 out of 10 scenes, particularly in CIDNet-9stg, achieving an average PSNR of 44.12dB and SSIM of 0.991. This significantly surpasses previous unfolding and end-to-end networks, and also in dual-camera CASSI systems, such as In2SET-9stg. A visual comparison is shown in Fig. 3. We provide the visual comparison of simulation Scene7 with 4 out of 28 spectral channels. In addition, we plot the spectral density curves of two regions in the top left RGB image. Our CIDNet-9stg achieves relatively higher spectral accuracy with reference spectra, demonstrating the effectiveness of our method. To verify the reconstructed quality of chromaticity and intensity, we compare the PSNR and SSIM of reconstructed chromaticity and intensity (with decomposition of HSIs), which is shown in Table 2. Our CIDNet achieves significant improvements in metrics of chromaticity and intensity. For dual-camera CASSI algorithm PIDS and In2SET, we achieve the best chromaticity metrics with a PSNR of 35.81 and a SSIM of 0.93. A visual comparison of reconstructed chromaticity is shown in Fig. 4. In this research, we used a real-world DCCHI measurement Ninja, taken from publicly available data as detailed in he2021fast . Fig. 6 illustrates the reconstruction results for four spectral bands in this scene, using two dual-camera CASSI reconstruction algorithms, PIDS and CID-TV, where CID-TV is the iterative CID algorithm using Total Variationa as Regularizer. The comparison highlights the superior image restoration quality of our model over other methods, validating its effectiveness and reliability in real-world applications.
@@ 3.2 Ablation Study
Table 4: Break-down ablation study on self-attention mechanism.
@@
Robustness of intensity-guided mask. We test the robustness of intensity-guided mask by employing this intensity mask to ADMM, DAUHST and MST, the result is shown in Tab.  5. We compare the reconstruction quality with/without intensity-guided mask and find that it is significantly improving the base (without intensity) in iterative, end-to-end and unfolding framework. Note that DAUHST achieves slightly higher metric than ours. However its flops and parameters are also greater.
Self-Attention Scheme Comparison. We compare Window-based Spectral Self-Attention (WSSA) zhang2024improving and TKSA and its variants within the encoder and decoder. We use $^{\prime}+^{\prime}$ to signify a parallel implementation of both attentions, and $^{\prime}-^{\prime}$ to represent an encoder-decoder implementation. Base-2 is CIDNet-3stg that removes the attention module. Tab. 4 shows the ablation results and LWSA-TKSA yields the most prominent improvement of 1.56dB PSNR compared with Base-1, which shows the effectiveness of our method.
@@ 4 Conclusion
We present CIDNet, a novel reconstruction framework for CASSI, which leverages a physically motivated chromaticity-intensity decomposition. By disentangling the hyperspectral image into a spatially smooth intensity map and a spectrally informative chromaticity cube, our method enables lighting-invariant reflectance modeling and better preserves spatial-spectral details. We have designed a hybrid spatial-spectral Transformer to recover the sparse and high-frequency chromaticity components and introduce a degradation-aware unfolding strategy with spatially adaptive noise modeling to handle anisotropic noise inherent in the dual-camera CASSI system. Extensive experiments on both simulated and real-world CASSI datasets validate the effectiveness of our approach, achieving state-of-the-art performance in spectral reconstruction and chromaticity fidelity. This work highlights the benefit of physics-aware decomposition and hybrid attention mechanisms in addressing the ill-posed inverse problem of CASSI reconstruction.
@@ 80. Guidelines:
Appendix A Intrinsic Modeling of Hyperspectral Radiance
@@ A.1 Multispectral Image Formation
Let $(u,v)$ be spatial coordinates and $\lambda\in[\lambda_{\min},\lambda_{\max}]$ the wavelength. Assume the scene reflectance is $R(u,v,\lambda)$, and the illumination spectral power is $L(\lambda)$, spatially uniform. Let $s(\lambda)$ be the spectral response function of the grayscale camera, which converts the spectral radiance to a scalar intensity. Then the intensity recorded by the grayscale camera at $(u,v)$ is:
$I(u,v)=\int_{\lambda_{\min}}^{\lambda_{\max}}s(\lambda)\cdot L(\lambda)\cdot R(u,v,\lambda)\,d\lambda.$ (25)
@@
$X(u,v,\lambda)=s(\lambda)\cdot L(\lambda)\cdot R(u,v,\lambda)$ (26)
@@
A.2 Illumination-Invariance of Chromaticity
@@
A.3 Application in Dual-Path CASSI Reconstruction
In a dual-camera system, the grayscale camera provides $I(u,v)$, representing spatial structure (intensity). The CASSI camera provides encoded spectral information. Our goal is to reconstruct the full spectral cube as:
$\mathbf{X}(u,v,\lambda)=I(u,v)\cdot\mathbf{C}(u,v,\lambda)$ (31)
That is, use the grayscale intensity as the **intensity prior**, and use inversion of the CASSI system to recover $\mathbf{X}(u,v,\lambda)$, which is lighting-invariant.
Appendix B Closed-form Solution of Data-fidelity Term
@@
Appendix C Traditional Optimization-based Methods
@@
Appendix D Additional Studies
D.1 Comparison of HSIs and chromaticity
@@ D.2 Noise map
D.3 Limitation and broader impact of our work

### v2

@@ 2.3 Hybrid Spatial-Spectral Transformer
We adopt a dual-branch design: the spatial attention branch captures intra-image textures through a Swin Transformer in Encoder, while the spectral attention branch is tailored to exploit sparse and locally correlated spectral features using a TopK spectral attention mechanism in Decoder. This asymmetric design is inspired by Zhang et al. (2024a) and motivated by experimental verification, where the asymmetric design has better reconstruction results.
@@
where $\mathcal{R}$ denotes the set of TopK ratios. This fusion across multiple sparsity levels enables the network to capture both dominant and complementary spectral correlations, further improving spectral detail preservation. and reshaped back to the original window structure. After processing all windows, the outputs are stitched to reconstruct the full feature map. This spectral TopK attention not only reduces the computational burden from dense $O(C^{2})$ to $O(KC)$, but also explicitly captures the structured sparsity observed in reflectance spectra—where only a few wavelengths contribute significantly. Empirically, we find that this strategy enhances spectral sharpness and suppresses irrelevant cross-band mixing. Following the standard Transformer architecture, we apply a conventional feedforward network after the sparse TopK attention, which is not the focus of our work.
@@
Inspired by recent advances in Swin Transformers Liu et al. (2021), we employ the Swin Transformer as our spatial modeling backbone. The input feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$ is divided into non-overlapping windows of size $M\times M$. Within each window, we perform multi-head self-attention (MSA) by computing
@@ 3 Experiments
Simulation Dataset. We adopt two widely used hyperspectral datasets: CAVE Yasuma et al. (2010) and KAIST Choi et al. (2017). The CAVE dataset contains 32 hyperspectral images with a spatial resolution of $512\times 512$. The KAIST dataset provides 30 high-resolution hyperspectral scenes of size $2704\times 3376$. We obtain ground-truth multi-spectral chromaticity and intensity using chromaticity-intensity decomposition Eq. (28). Following prior works Meng et al. (2020b); Miao et al. (2019); Cai et al. (2022a), we use all CAVE images for training and select 10 scenes from KAIST for evaluation.
Implementation Details. Our model is implemented in PyTorch and trained using the Adam optimizer Kingma (2014) for 300 epochs. The initial learning rate is set to $4\times 10^{-4}$ and updated using a cosine annealing schedule. We employ the $\ell_{2}$ loss between the reconstructed chromaticity and ground-truth chromaticity as the objective function. For training, we randomly extract 3D hyperspectral patches from each scene. For simulated data, the patch size is $256\times 256\times 28$, for real-world data, we use patches of size $350\times 260\times 26$. In simulation experiments, the forward imaging model is configured with a dispersion shift step $d=2$, directing dispersion along the horizontal axis (rightward). In real-world scenarios, we assume a vertical dispersion direction and set $d=1$, consistent with the dual-camera hardware setup. All experiments are conducted in Nvidia A40 GPU.
@@
Comparing Methods. We compared the HSIs and chromaticity reconstruction performance of our CIDNet with other 7 SOTA methods, including DeSCILiu et al. (2018), GAP-NetMeng et al. (2020a), MST-LCai et al. (2022a), DAUHST-9stgCai et al. (2022b), SSR-9stgZhang et al. (2024b) and two dual-camera CASSI algorithm: PIDSChen et al. (2023) and In2SET-9stgWang et al. (2024). PIDS is compared with RGB image as guidance (in original paper). We evaluate the reconstruction performance from two perspectives: hyperspectral images (HSIs) and chromaticity. Unlike the baseline methods that directly reconstruct HSIs, our method assumes that the intensity is known and focuses on reconstructing the chromaticity. For a fair comparison in the HSI domain, we obtain our reconstructed HSIs by multiplying the recovered chromaticity with the known intensity. Conversely, for chromaticity-level comparison, we perform chromaticity–intensity decomposition on the HSIs reconstructed by the baseline methods to extract their chromaticity components. The reconstruction quality of HSIs is evaluated using peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM).
@@ 3.1 Quantitative Results
As shown in Tab. 2, we compare the PSNR and SSIM of HSIs with SOTA methods. our proposed CIDNet excels in 8 out of 10 scenes, particularly in CIDNet-9stg, achieving an average PSNR of 44.12dB and SSIM of 0.991. This significantly surpasses previous unfolding and end-to-end networks, and also in dual-camera CASSI systems, such as In2SET-9stg. A visual comparison is shown in Fig. 3. We provide the visual comparison of simulation Scene7 with 4 out of 28 spectral channels. In addition, we plot the spectral density curves of two regions in the top left RGB image. Our CIDNet-9stg achieves relatively higher spectral accuracy with reference spectra, demonstrating the effectiveness of our method. To verify the reconstructed quality of chromaticity and intensity, we compare the PSNR and SSIM of reconstructed chromaticity and intensity (with decomposition of HSIs), which is shown in Table 3. Note that our method assumes a known intensity. Therefore, to ensure a fair comparison of intensity, we decompose the reconstructed HSIs to extract their intensity component for evaluation. Our CIDNet achieves significant improvements in metrics of chromaticity and intensity. For dual-camera CASSI algorithm PIDS and In2SET, we achieve the best chromaticity metrics with a PSNR of 35.81 and a SSIM of 0.93. A visual comparison of reconstructed chromaticity is shown in Fig. 4. In this research, we used a real-world DCCHI measurement Ninja, taken from publicly available data as detailed in He et al. (2021). Fig. 6 illustrates the reconstruction results for four spectral bands in this scene, using two dual-camera CASSI reconstruction algorithms, PIDS and CID-TV, where CID-TV is the iterative CID algorithm using Total Variationa as Regularizer, details can be found in supplement materials. The comparison highlights the superior image restoration quality of our model over other methods, validating its effectiveness and reliability in real-world applications.
@@ 3.2 Ablation Study
Table 5: Break-down ablation study on spectral self-attention mechanism.
@@
Robustness of intensity-guided mask. We test the robustness of intensity-guided mask $\mathbf{M}^{\prime}$ by employing this intensity mask to ADMM, DAUHST and MST, the result is shown in Tab.  6. We compare the reconstruction quality with/without intensity-guided mask and find that it is significantly improving the base (without intensity) in iterative, end-to-end and unfolding framework. Note that DAUHST achieves slightly higher metric than ours. However its flops and parameters are also greater.
Self-attention scheme comparison. We compare Window-based Spectral Self-Attention (WSSA) Zhang et al. (2024b) and spectral TKSA and its variants within the encoder and decoder. We use $^{\prime}+^{\prime}$ to signify a parallel implementation of both attentions (each processing half of the feature channels), and $^{\prime}-^{\prime}$ to represent an encoder-decoder implementation. Base-2 is CIDNet-3stg that removes the attention module. Tab. 5 shows the ablation results and LWSA-TKSA yields the most prominent improvement of 1.56dB PSNR compared with Base-2, which shows the effectiveness of our method.
@@ 4 Conclusion
We present CIDNet, a novel reconstruction framework for CASSI, which leverages a physically motivated chromaticity-intensity decomposition. By disentangling the hyperspectral image into a spatially smooth intensity map and a spectrally informative chromaticity cube, our method enables lighting-invariant reflectance modeling and better preserves spatial-spectral details. We have designed a hybrid spatial-spectral Transformer to recover the sparse and high-frequency chromaticity components and introduced a degradation-aware unfolding strategy with spatially adaptive noise modeling to handle anisotropic noise inherent in the dual-camera CASSI system. Extensive experiments on both simulated and real-world CASSI datasets validate the effectiveness of our approach, achieving state-of-the-art performance in spectral reconstruction and chromaticity fidelity. This work highlights the benefit of physics-aware decomposition and hybrid attention mechanisms in addressing the ill-posed inverse problem of CASSI reconstruction.
Acknowledgments
This work was supported by National Key R&D Program of China (2024YFF0505603), the National Natural Science Foundation of China (grant number 62271414), Zhejiang Provincial Distinguished Young Scientist Foundation (grant number LR23F010001), Zhejiang ‘Pioneer’ and ‘Leading Goose’ R&D Program (grant number 2024SDXHDX0006, 2024C03182), the Key Project of Westlake Institute for Optoelectronics (grant number 2023GD007), the 2023 International Sci-tech Cooperation Projects under the purview of the ‘Innovation Yongjiang 2035’ Key R&D Program (grant number 2024Z126), National Natural Science Foundation of China (grant number 62501487). (Corresponding author: Xin Yuan.)
@@ 80. Guidelines:
Appendix A Appendix
In the supplementary material, we provide more details that are not in out main paper:
(a) Multispectral Image Formation in Sec. A.1.
(b) Proof of Illumination-Invariance of Chromaticity in Sec. A.2.
(c) Proof of PAN-Intensity Equivalence in Sec. A.3. We demonstrate the PAN image in dual-camera CASSI is equivalent to the Intensity image.
(d) Closed-form Solution of Data-fidelity Term in Sec. A.4.
(e) Traditional Optimization-based Methods in Sec. A.5. Some results are visualized.
(f) Visual Comparison of HSIs and Chromaticity in Sec. A.6.
(g) Noise Map Visualization of DNEM in Sec. A.7.
(h) Limitation and Broader Impact of Our Work in Sec. A.8
@@ A.1 Multispectral Image Formation
Let $(u,v)$ be spatial coordinates and $\lambda\in[\lambda_{\min},\lambda_{\max}]$ the wavelength. Assume the scene reflectance is $\mathbf{R}(u,v,\lambda)$(this is not the same as data-prior term), and the illumination spectral power is $L(\lambda)$, spatially uniform. Let $s(\lambda)$ be the spectral response function of the grayscale camera, which converts the spectral radiance to a scalar intensity. Then the intensity recorded by the grayscale camera at $(u,v)$ is:
$\mathbf{I}(u,v)=\int_{\lambda_{\min}}^{\lambda_{\max}}s(\lambda)\cdot L(\lambda)\cdot\mathbf{R}(u,v,\lambda)\,d\lambda.$ (25)
@@
$\mathbf{X}(u,v,\lambda)=s(\lambda)\cdot L(\lambda)\cdot\mathbf{R}(u,v,\lambda)$ (26)
@@
A.2 Proof of Illumination-Invariance of Chromaticity
@@
A.3 Proof of PAN-Intensity Equivalence
Proposition A.1 (PAN-Intensity Equivalence Under Uniform Illumination).
In a dual-camera system comprising a CASSI sensor and a grayscale PAN camera exposed under the same illumination $L(\lambda)$, let $s(\lambda)$ denote the spectral response of the camera. The PAN image $\mathbf{I}_{\mathrm{PAN}}(u,v)$ provides a relative estimate of the scene intensity $\mathbf{I}(u,v)$ defined by the chromaticity-intensity decomposition of the hyperspectral image $\mathbf{X}(u,v,\lambda)$. That is,
$\mathbf{I}_{\mathrm{PAN}}(u,v)\approx k\cdot\mathbf{I}(u,v),$ (31)
where $k$ is a scalar constant that is approximately invariant across spatial coordinates $(u,v)$.
Proof.
Our derivation begins with the Retinex theory, which decomposes an image into chromaticity and intensity components. For hyperspectral images, this decomposition is generalized as:
$\mathbf{X}(u,v,\lambda)=\mathbf{C}(u,v,\lambda)\cdot\mathbf{I}(u,v),$ (32)
where the chromaticity and intensity are defined as:
$\mathbf{C}(u,v,\lambda)$ $=\frac{\mathbf{X}(u,v,\lambda)}{\int\mathbf{X}(u,v,\lambda^{\prime})d\lambda^{\prime}},$ (33)
$\mathbf{I}(u,v)$ $=\int\mathbf{X}(u,v,\lambda)d\lambda.$ (34)
In our dual-camera setup with CASSI and PAN sensors under identical illumination $L(\lambda)$, the PAN image formation is modeled as:
$\mathbf{I}_{\mathrm{PAN}}(u,v)=\int s(\lambda)\cdot L(\lambda)\cdot\mathbf{X}(u,v,\lambda)d\lambda.$ (35)
Substituting the $\mathbf{X}$ into the product of chromaticity and intensity yields:
$\mathbf{I}_{\mathrm{PAN}}(u,v)=\int s(\lambda)L(\lambda)[\mathbf{C}(u,v,\lambda)\cdot\mathbf{I}(u,v)]d\lambda.$ (36)
Since $\mathbf{I}(u,v)$ is independent of wavelength $\lambda$, it can be factored out of the integral:
$\mathbf{I}_{\mathrm{PAN}}(u,v)=\mathbf{I}(u,v)\cdot\int s(\lambda)L(\lambda)\mathbf{C}(u,v,\lambda)d\lambda.$ (37)
The key insight is that $\mathbf{C}(u,v,\lambda)$ is normalized ($\int\mathbf{C}(u,v,\lambda^{\prime})d\lambda^{\prime}=1$) and exhibits smooth spectral variation, while $s(\lambda)L(\lambda)$ acts as a broadband low-pass filter. This justifies the approximation:
$\int s(\lambda)L(\lambda)\mathbf{C}(u,v,\lambda)d\lambda\approx k,$ (38)
where $k$ is a spatial-invariant scalar constant. Therefore, we obtain:
$\mathbf{I}_{\mathrm{PAN}}(u,v)\approx k\cdot\mathbf{I}(u,v).$ (39)
To address scale ambiguity, we normalize the PAN image to $[0,1]$ during training and inference, which justifies using PAN as a relative intensity estimate in our chromaticity-intensity framework. ∎
A.4 Closed-form Solution of Data-fidelity Term
@@
A.5 Traditional Optimization-based Methods
@@
A.6 Visual Comparison of HSIs and Chromaticity
@@ A.7 Noise Map
A.8 Limitation and Broader Impact of Our Work